How much force is needed to punch a circular hole 1.2cm in radius in a metal plate of thickness 0.3 cm? The shear modulus of the metal is 150.0 x 106 N·m

1
Expert's answer
2022-05-24T13:56:13-0400
"F=SA=S*\pi *d*t\\
=150*10^6*3.14*(2*0.012)*0.003\\
=34000\:\rm N"